The ratio of tweedle-dees to tweedle-dums was 2 to 7.if a total of 108 were present,how many of them were tweedle-dums
Answer by checkley77(12844)

Ratio 2 to 7. total of 108.
2n=7n=9n
9n=108
n=108/9
n=12 is the ratio multiplier.
2*12=24
7*12=84 Tweedle-dums.
Proof:
24/84=2/7
24+84=108